OpenStudy (anonymous):

pls help medal and fan 1 over 6 multiplied by negative 1 over 2.

OpenStudy (anonymous):

Answer: = 1/6 × -1/2 = (1 × -1) / (6 × 2) = -1/12 = -1/12 Btw you should use a calculator sometime.

OpenStudy (anonymous):

wouldn't that be -1/12?

OpenStudy (anonymous):

thankss